# Prototype Augmented Hypernetworks for Continual Learning

### v1

@@ 2 Related Work
Our work intersects hypernetworks and prototype‑based methods for continual learning. Hypernetworks [7] generate parameters for a target network and have been adapted to CL by conditioning on task identity to produce either full‑network weights or selected layers [25, 8]. PAH keeps the same generator idea but feeds it learnable task prototypes to dynamically generate weights only for the task-specific classifier head, avoiding the need to store separate heads The prototypes-compact class exemplars-have a long history in few‑shot learning [22] and as anchors for replay or nearest‑neighbour decisions in CL [20, 9]. In PAH prototypes act instead as a task embedding that conditions the hypernetwork; while knowledge distillation [15] limits forgetting. The approach differs from Prototype Augmentation and Self‑Supervision (PASS) [28], which keeps one prototype per class and feeds those fixed vectors directly to the classifier. $PAH$ stores no classifier heads and never sends prototypes directly to softmax at test time.
Figure 1: Overview of the Prototype-Augmented Hypernetworks ($PAH$) architecture at task t. The hypernetwork generates task-specific classifier weights conditioned on learnable task prototypes $e_{t}$, which encode task-specific information. This design eliminates the need to store separate classifier heads for each task and effectively addresses catastrophic forgetting through knowledge distillation.
@@ 4 Methodology
These prototypes provide the hypernetwork with a compact yet rich encoding, enabling it to generate more precise task-specific classifier weights. Unlike fixed embeddings, prototypes are dynamically optimized during training, ensuring alignment with evolving backbone representations. This structured approach allows the hypernet to leverage class-specific information, improving adaptability while reducing the need to store separate classifier heads.
@@ 6 Results
On TinyImageNet, $PAH$ achieves 63.65% accuracy and just 4.43% forgetting. In contrast, replay-based methods such as GCR [23] and DER [1] suffer from around 40% forgetting despite storing samples. Similarly, hypernetwork-based methods HN-2 and HN-3 [8] lag behind $PAH$ in both metrics. These results show the effectiveness of combining hypernets with prototypes to achieve minimal forgetting and strong accuracy, establishing $PAH$ as a new benchmark in CL.

### v2

@@ 2 Related Work
Our work intersects hypernetworks and prototype‑based methods for continual learning. Hypernetworks [7] generate parameters for a target network and have been adapted to CL by conditioning on task identity to produce either full‑network weights or selected layers [25, 8]. PAH keeps the same generator idea but feeds it learnable task prototypes to dynamically generate weights only for the task-specific classifier head, avoiding the need to store separate heads. The prototypes–compact class exemplars–have a long history in few‑shot learning [22] and as anchors for replay or nearest‑neighbour decisions in CL [20, 9]. In PAH prototypes act instead as a task embedding that conditions the hypernetwork; while knowledge distillation [15] limits forgetting. The approach differs from Prototype Augmentation and Self‑Supervision (PASS) [28], which keeps one prototype per class and feeds those fixed vectors directly to the classifier. $PAH$ stores no classifier heads and never sends prototypes directly to softmax at test time.
Figure 1: Overview of the Prototype-Augmented Hypernetworks ($PAH$) architecture at task k. The hypernetwork generates task-specific classifier weights conditioned on learnable task prototypes $e_{k}$, which encode task-specific information. This design eliminates the need to store separate classifier heads for each task and effectively addresses catastrophic forgetting through knowledge distillation.
@@ 4 Methodology
These prototypes provide the hypernetwork with a compact yet rich encoding, enabling it to generate more precise task-specific classifier weights. Unlike fixed embeddings, prototypes are dynamically optimized during training, learning to capture class-specific feature distributions, and ensuring alignment with evolving backbone representations. This approach allows the hypernet to leverage task- and class-specific information, improving adaptability while reducing the need to store separate classifier heads.
@@ 6 Results
On TinyImageNet, $PAH$ achieves 63.65% accuracy and just 4.43% forgetting. In contrast, replay-based methods such as GCR [23] and DER [1] suffer from around 40% forgetting despite storing samples. Similarly, hypernetwork-based methods HN-2 and HN-3 [8] lag behind $PAH$ in both metrics. These results show the effectiveness of combining hypernets with prototypes to achieve minimal forgetting and strong accuracy, establishing $PAH$ as a new benchmark in CL. This low classifier-level forgetting suggests PAH effectively mitigates the representation drift impact where it is most damaging.